# Learning Policies for Dynamic Coalition Formation in Multi-Robot Task Allocation

### v1

@@ V-B Results
IV -
Ours -
We conducted a series of simulations to evaluate the proposed framework and the baseline across various environment settings, characterized by different task spawn configurations and task distributions, while keeping other parameters constant. A summary of all parameters is provided in Table III(b). In the homogeneous task distribution scenario, the task spawning area is four times larger than in the non-homogeneous case. To maintain consistent robot and task densities across both distributions, the number of robots and tasks is proportionally increased by a factor of four in the homogeneous case. For this analysis, we specifically focused on the M2 and 4xM2 settings, as detailed in Table II.
Fig. 3 compares the performance of our model, the baseline (PCFA), and various ablation studies on the components of our model for both non-homogeneous and homogeneous task distributions. Table III(c) presents an overview of the models examined in this simulation, highlighting their components and referencing relevant existing works associated with these models. Each model was trained in a setting-specific manner, spanning five task spawn configurations (Bernoulli spawn with probabilities $p=0.01,0.02,0.03,0.04$ and instant respawn) for each task distribution.
Models I, II, and III in this study are relevant to existing works in the literature ([34], [17], and [23], respectively), but with notable distinctions: (i) all three models are trained using MAPPO (CTDE) within a Dec-POMDP framework, unlike other works that rely on the MDP assumption with individual rewards; (ii) while [17] focuses on individual-robot training scenarios, Model II is trained for MRTA; (iii) in [23], only level-1 tasks are considered, and an algorithm similar to Independent Q-Learning (IQL) [42], which encourages competition rather than cooperation, is used instead of MAPPO.
Our method consistently outperforms PCFA in all configurations. Notably, as the spawn rate $p$ increases, PCFA’s performance declines, suggesting that it struggles to coordinate robots to prioritize and allocate tasks effectively across varying levels of complexity, despite its use of multiple communication rounds to reach consensus at each step.
@@
Figure 4: The performance of the proposed framework is analyzed in relation to the average task revision rate, with data points collected over the last 500,000 training steps for each task spawn scenario. As shown, the robots gradually learn an optimal task revision strategy during training, resulting in enhanced overall performance.
@@
We observe that task revision significantly enhances performance, particularly in highly dynamic settings such as instant respawn and Bernoulli spawn with $p=0.04$. These findings align with our expectation that, in environments where tasks spawn frequently and dynamically, robots must plan for long-term task execution while continuously reassessing and updating their plans. While frequent adjustments to task execution plans might intuitively seem disruptive to coalition formation among multiple robots working on high-level tasks, our proposed method enables robots to learn an optimal task revision strategy, ultimately leading to superior performance. Fig. 4 further demonstrates how our model effectively learns this strategy across various task spawn settings.
@@
We also conducted ablation studies to evaluate the impact of intention sharing and task revision on our model’s performance. Models without task revision are limited to selecting a new task only upon reaching assigned tasks, regardless of whether it executes the tasks or not. As shown in Fig. 3, Model II (without intention sharing or task revision) underperforms both Model III (with intention sharing only) and Model IV (with task revision only), highlighting the importance of learning the optimal timing for task revision and its communication with neighboring robots. The performance difference between Models III and IV suggests that, in task allocation scenarios requiring coalition formation, determining the optimal timing for task revision is essential.
@@
In the case of a homogeneous task distribution, Model I performs comparably to our model across all Bernoulli spawn settings but underperforms in the instant respawn scenario. This outcome suggests that homogeneous task scenarios with Bernoulli spawn do not necessitate long-term planning. However, in the instant respawn scenario, the constant task density discourages simplistic approaches like those employed by Model I, resulting in its reduced performance.
@@ V-C Scalability and Generalizability Evaluation
To evaluate scalability, we first trained policies in a homogeneous task distribution environment with a fixed number of robots. The trained policy was then evaluated in environments with progressively larger numbers of robots, while proportionally scaling the environment size $W$ and the number of tasks to maintain constant robot density $(\lambda_{N}=N/{W^{2}})$ and task density $(\lambda_{T,t}=T_{t}/{W^{2}})$, where $T_{t}$ represents the number of tasks available at time $t$. The training environment settings were consistent with the homogeneous task simulation but utilized a smaller grid size ($W=10$). Only the instant respawn scenario (where $T_{t}$ remains constant over $t$) was explored, using the M2 task setting (see Table II), and multiple robot densities were considered: $\lambda_{N}\in\{0.05,0.1,0.15,0.2\}$.
Fig. 5 illustrates the performance of trained policies as a function of $N\in[10,1000]$. We observe that the policy’s performance scales linearly with $N$ for all robot densities evaluated. These findings validate the scalability of our method, demonstrating its effectiveness in managing larger numbers of robots for task allocation within proportionally scaled environments.
Figure 5: Our model’s performance vs. number of robots.
Figure 6: Generalizability results.
To evaluate the generalizability of our method across environments with varying task settings, we train and evaluate our model on a diverse set of environments, each characterized by different task distributions and levels. As detailed in Table II, we designed nine task settings, categorized as easy, medium, or hard, each defined by a specific combination of task levels and quantities.
@@
Fig. 6 presents the generalizability scores of policies trained in easy, medium, and hard task settings, as well as in a setting combining all three. We observe that the policy trained across all settings can generalize better and more consistently. This indicates that this policy performs close to the highest performance that setting-specific policies can attain. Based on these results, we conclude that our method enables trained policies to allow robots to achieve near-optimal performance across environments with different task level distributions when trained in diverse environment settings.
@@ VI Conclusions
In this work, we introduced a learning-based framework to address decentralized dynamic coalition formation. By leveraging the MAPPO algorithm and incorporating spatial action maps, motion control, task allocation revision, and intention sharing, our model demonstrated significant performance improvements over existing approaches. Extensive simulation studies highlighted the scalability of our framework with an increasing number of robots and its generalizability, performing well across diverse environments.
Future directions include enhancing the learning algorithm by integrating learned communication strategies (instead of predefined intention-based communication) and incorporating mechanisms for achieving consensus among robots, particularly for bandwidth-abundant environments.
@@
We implemented PCFA and integrated it with GyMRTA for the performance comparison. To ensure a fair comparison with our method, the following modifications are applied:
  • Partial Observability: In PCFA, the utility function for a task is set to zero if the task lies outside the robot’s view range $R_{\text{view}}$, preventing robots from forming coalitions for tasks they cannot observe. Additionally, the communication graph is updated at each time step, with links existing only between robots that are within the communication range $R_{\text{comm}}$.
  • Task Assignment Revision: To align with our method, robots are allowed to revise their task assignments at every time step. This is achieved by recomputing all four stages of PCFA at every time step. While this modification increases communication overhead compared to the original PCFA, it enables a fair comparison.
The utility function was adjusted to align with the reward function used in our framework:
$s_{k}(t_{\text{ETA}})=l_{k}^{2}\,e^{-2(t_{\text{ETA}}+t_{w}-t_{k}^{0})},$
where $l_{k}$ is the task level, $t_{\text{ETA}}$ is the estimated time of arrival, $t_{w}$ is the wait time, and $t_{k}^{0}$ is the task’s initial spawn time.
The modified PCFA implementation is available in our code repository.

### v2

@@ IV-B1 Multi-Agent Proximal Policy Optimization (MAPPO)
Ours
IV-B2 Learning Task Allocation and Revision
To address this challenge, we extend the MAPPO framework to train a policy capable of both task allocation and revision. Unlike existing approaches [23, 22], where task revisions occur only after a robot reaches its assigned task location and either completes it or fails due to the absence of a coalition, our trained policy proactively reallocates tasks as needed, based on evolving local observations and information shared by neighboring robots. As demonstrated in §V, this integrated approach enables our framework to concurrently learn task allocation and revision, ensuring optimization of the overall objective in the Dec-POMDP environment.
V Evaluation
We conducted a series of simulations to evaluate: (i) the performance of our proposed framework compared to existing methods, including the state-of-the-art market-based approach PCFA, [^3] The original PCFA [8] was modified to accommodate robots with limited view and communication ranges and to enable task reassignment. Details of these modifications are provided in the appendix. (ii) its scalability with respect to the number of robots, and (iii) its generalizability across a range of task allocation scenarios with varying task difficulties. Furthermore, we present an ablation study to underscore the critical role of task revision in dynamic coalition formation.
V-A Simulation Setup and Performance Metric
To evaluate our proposed method, we developed GyMRTA, a Gym environment for discrete-space, discrete-time MRTA, and extended the existing MAPPO implementation from the EPyMARL [36] package. Our extensions include features such as task revision, action filtering, multi-dimensional observation and action spaces, support for custom state space, and tailored actor and critic Neural Network (NN) architectures. The hyperparameters of MAPPO, presented in Table II(a), remain constant across all simulations. Each episode is limited to 100 time steps ($T=99$) to ensure a finite episodic reward.
At the start of each episode, the initial positions of robots and tasks are randomly sampled. Table III details on the initial number of tasks per level across various task settings used in the simulations. The environment is categorized based on task distribution as either homogeneous or non-homogeneous.
TABLE III: Task settings.
Setting Name Tasks Setting Name Tasks
1 2 3 1 2 3
Easy Hard
E1 15 - - H1 - 5 5
E2 10 - - H2 - - 10
E3 5 5 - H3 - - 5
Medium Non-Homogeneous
M1 6 4 4 M2 4 3 3
M2 4 3 3 Homogeneous
M3 2 2 1 4$\times$×M2 16 12 12
In the homogeneous scenario, tasks are evenly distributed across the grid, maintaining a uniform density. Conversely, the non-homogeneous scenario features an uneven distribution of tasks, leading to variations in task density. To define non-homogeneous configurations, we adopt the random corner patch distribution: the grid is divided into four equal-sized corner patches, and at the start of each episode, one patch is randomly chosen to contain all tasks. Within the chosen patch, the task density is uniform, while the rest of the grid remains empty. The non-homogeneous scenarios are used to evaluate whether the learned policy enables robots to form effective coalitions for task execution while efficiently locating tasks.
V-B Simulation Results
(a) Non-homogeneous tasks.
(b) Homogeneous tasks.
@@ V-B Simulation Results
Figure 4: The performance of the proposed framework is analyzed with respect to the task revision rate, i.e., the average number time steps that each robot changes its task selection, with data points collected over the last 500,000 training steps for each task spawn scenario.
We conducted a series of simulations to evaluate the proposed framework and the baseline across various environment settings, characterized by different task spawn configurations and distributions, while keeping other parameters constant. A summary of all simulation parameters is provided in Table II(b). In the homogeneous task distribution scenario, the task spawning area is four times larger than in the non-homogeneous case. To ensure consistent robot and task densities across both distributions, the number of robots and tasks in the homogeneous scenario is proportionally increased by a factor of four. For this purpose, we use the M2 and 4$\times$×M2 settings for the non-homogeneous and homogeneous scenarios, respectively, as described in Table III.
Fig. 3 compares the performance of our model, the baseline (PCFA), and various ablation studies on the components of our model across non-homogeneous and homogeneous task distributions and five task spawn configurations (Bernoulli spawn with probabilities $p=0.01,0.02,0.03,0.04$ and instant respawn), spanning 10 different settings. Table II(c) provides details of the models used in the ablation studies, highlighting their components and referencing relevant existing works associated with each model. [^4] Spatial action refers to defining the action in Dec-POMDP as the location of a task. Consequently, in model I, which does not incorporate spatial action, an action is defined as a neighboring cell on the grid that a robot can move to within a single time step. Each model was trained and evaluated in every environment setting.
Models I, II, and III in this study are closely related to existing works [35], [17], and [23], respectively, but include the following necessary modifications: (i) All three models are trained using MAPPO (CTDE) within a Dec-POMDP framework, whereas the original references rely on the MDP assumption with individual rewards. (ii) While [17] focuses on individual-robot training scenarios, model II is trained for MRTA using our framework. (iii) In [23], only level-1 tasks are considered, and a method similar to Independent Q-Learning (IQL) [42], which encourages competition rather than cooperation among multiple robots, is used, while model III uses MAPPO.
As shown in Fig. 3, our method consistently outperforms PCFA in all configurations. Notably, as the spawn rate $p$ increases, PCFA’s performance declines, indicating its difficulty in coordinating robots to effectively prioritize and allocate tasks of varying difficulty levels. Moreover, we observe that task revision significantly improves performance, particularly in highly dynamic scenarios such as the instant respawn and Bernoulli spawn with $p=0.04$. These results align with our expectation that in environments where tasks spawn frequently and dynamically, robots must continuously reassess and update their long-term task execution plans.
@@
While frequent adjustments to task execution plans might seem disruptive to coalition formation among multiple robots working on high-level tasks, our proposed method enables robots to learn an optimal task revision strategy, ultimately resulting in superior performance. Figure 4 illustrates how our model effectively learns this strategy across various task spawn configurations. As shown, the performance of our framework tends to decline with excessive task assignment revisions. Conversely, without task revision, – corresponding to model III in the ablation study – the performance is significantly suboptimal, as illustrated in Fig. 3.
@@
The ablation studies evaluate the impact of intention sharing and task revision on our model’s performance. Models without task revision are limited to selecting a new task only upon reaching assigned tasks, regardless of whether it executes the tasks or not. As shown in Fig. 3, model II (without intention sharing or task revision) underperforms both model III (with intention sharing only) and our model (with both intention sharing and task revision), highlighting the importance of learning the optimal timing for task revision and its communication with neighboring robots.
@@
In the case of a homogeneous task distribution, model I performs comparably to our model across all Bernoulli spawn settings but underperforms in the instant respawn scenario. This outcome indicates that homogeneous task scenarios with Bernoulli spawn do not necessitate long-term planning. However, in the instant respawn scenario, the constant task density discourages simplistic approaches like those employed by model I, resulting in its reduced performance.
@@ V-C Scalability and Generalizability Evaluation
To evaluate scalability, we first trained policies in a homogeneous task distribution environment ($W=10$) with a fixed number of robots. The trained policy was then tested while progressively scaling the environment size $W$ and the number of tasks $T_{t}$ to maintain constant robot density $(\lambda_{N}=N/{W^{2}})$ and constant task density $(\lambda_{T_{t}}=T_{t}/{W^{2}}=0.1)$. To ensure $T_{t}$ is time-invariant, we focused on the instant respawn scenario, using the M2 task setting (see Table III). Multiple robot densities were evaluated: $\lambda_{N}\in\{0.05,0.1,0.15,0.2\}$.
Fig. 5(a) illustrates the performance of trained policies as a function of $N\in[10,1000]$. We observe that the policy’s performance scales linearly with $N$ for all robot densities evaluated. These results validate the scalability of our method, demonstrating its effectiveness in managing larger numbers of robots for task allocation while maintaining constant robot and task densities.
Figure 5: Scalability (a) and generalizability (b) results.
To evaluate the generalizability of our method across environments with varying task settings, we train and evaluate our model on a diverse set of environments, each characterized by different task distributions and levels. As detailed in Table III, we designed nine task settings, categorized as easy, medium, or hard, each defined by a specific combination of task levels and quantities.
@@
Fig. 5(b) shows the generalizability scores of policies trained in easy, medium, and hard task settings, as well as in a combined setting encompassing all three. The results demonstrate that the policy trained across all settings achieves better and more consistent generalization. Notably, this policy performs close to the highest performance attainable by setting-specific policies. These results indicate that our method enables trained policies to allow robots to achieve near-optimal performance across environments with varying task-level distributions when trained in diverse environment settings.
@@ VI Conclusions
We introduced a learning-based framework to address decentralized dynamic coalition formation. By leveraging the MAPPO algorithm and incorporating spatial action maps, motion planning, task allocation revision, and intention sharing, our model demonstrated significant performance improvements over existing approaches. Extensive simulation studies highlighted the scalability of our framework with an increasing number of robots and its generalizability, performing well across diverse environments.
Future directions include improving the learning algorithm by integrating learned communication strategies in place of predefined intention-based communication. Additionally, we aim to explore mechanisms for achieving consensus among robots to expedite coalition formation, particularly for bandwidth-abundant environments.
@@
We implemented PCFA and integrated it with GyMRTA for the performance comparison. To ensure a fair comparison with our method, the following modifications are applied: (i) Utility Function: The utility function $s_{k}$, associated with the $k$-th task, was adjusted to align with the reward function used in our framework: $s_{k}(t_{\text{ETA}})=l_{k}^{2}\,e^{-2(t_{\text{ETA}}+t_{w}-t_{k}^{0})},$ where $l_{k}$ is the task level, $t_{\text{ETA}}$ is the estimated time of arrival, $t_{w}$ is the wait time, and $t_{k}^{0}$ is the task’s initial spawn time. (ii) Partial Observability: In PCFA, the utility function for a task is set to zero if the task lies outside the robot’s view range $R_{\text{view}}$, preventing robots from forming coalitions for tasks they cannot observe. Additionally, the communication graph is updated at each time step, with links existing only between robots that are within the communication range $R_{\text{comm}}$. (iii) Task Assignment Revision: To align with our method, robots are allowed to revise their task assignments at every time step. This is achieved by recomputing all four stages of PCFA at every time step. While this modification increases communication overhead compared to the original PCFA, it allows a fair comparison in our simulation study.